An investor invested a total of $2500 in two mutual funds. One fund earned a 8% profit while the other earned a 3% profit. If the investors total profit was $110, how much was invested in each mutual fund?
Answer by mananth(16949)

Mutual fund I 8.00% per annum ------------- Amount invested =x
Mutual fund II 3.00% per annum ------------ Amount invested = y
2500
Interest----- 110

Mutual fund I 8.00% per annum ---x
Mutual fund II 3.00% per annum ---y
Sum of investments
x + 1 y= 2500 -------------1
Interest earned on both investments
8.00% x + 3.00% y= 110
Multiply by 100
8 x + 3 y= 11000.00 --------2
Multiply (1) by -8
we get
-8 x -8 y= -20000.00
Add this to (2)
0 x -5 y= -9000
divide by -5
y = 1800
Mutual fund I 8.00% $ 700
Mutual fund II 3.00% $ 1800

CHECK
1800 --------- 3.00% ------- 54.00
700 ------------- 8.00% ------- 56.00
Total -------------------- 110.00

m.ananth@hotmail.ca